# HAMLET: Hyperadaptive Agent-based Modeling for Live Embodied Theatrics

### v1

@@ Abstract
Creating an immersive and interactive theatrical experience is a long-term goal in the field of interactive narrative. The emergence of large language model (LLM) is providing a new path to achieve this goal. However, existing LLM-based drama generation methods often result in AI agents that lack initiative and cannot interact with the physical environment. Furthermore, these methods typically require detailed user input to drive the drama. These limitations reduce the interactivity and immersion of online real-time performance. To address the above challenges, we propose HAMLET, a multi-agent framework focused on drama creation and online performance. Given a simple topic, the framework generates a narrative blueprint, guiding the subsequent improvisational performance. During the online performance, each actor is given an autonomous mind. This means that actors can make independent decisions based on their own background, goals, and emotional state. In addition to conversations with other actors, their decisions can also change the state of scene props through actions such as opening a letter or picking up a weapon. The change is then broadcast to other related actors, updating what they know and care about, which in turn influences their next action. To evaluate the quality of drama performance, we designed an evaluation method to assess three primary aspects, including character performance, narrative quality, and interaction experience. The experimental evaluation shows that HAMLET can create expressive and coherent theatrical experiences. Our code, dataset and models are available at https://github.com/HAMLET-2025/HAMLET.
Figure 1: The HAMLET framework creates AI drama in two main stages. First, during offline planning, a collaborative workflow of agents including the actor designer, plot designer, and reviewer creates initial materials, which are then integrated by a director agent into a structured narrative blueprint. This blueprint then guides the subsequent online performance, where a control system composed of a planner, transfer, and advancer directs a dynamic and improvisational theatrical experience. The figure illustrates this entire process using the play-within-a-play scene from Hamlet, Act III, Scene II.
Introduction
@@
In recent years, large language models (LLMs) have demonstrated strong ability in various generative tasks, especially in creative fields such as story creation (Li et al. 2024) and role-playing (Tu et al. 2024). These models are able to generate fluent and imaginative texts, providing a foundation for the development of interactive narrative. When applying these capabilities to the creation of drama, the exploration can be divided into two parts: drama generation and drama performance.
@@
Existing studies have explored drama generation using language models, employing methods such as hierarchical approaches (Fan, Lewis, and Dauphin 2018; Yao et al. 2019) and the modular generation of script structures (Mirowski et al. 2023). However, significant challenges exist in applying LLMs to structured drama generation, especially in scenarios requiring online real-time enactment. In interactive drama performance, LLM agents take on the role of performing the narrative. In contrast to free-form story creation, drama requires each actor to make decisions and take actions consistent with their profiles and the overall plot progression. These interactions must unfold across a series of scenes to collaboratively advance the narrative. Furthermore, existing methods typically require detailed user input, such as a full story outline (Wu et al. 2024) or elaborate guiding paragraphs (Wu et al. 2025c), which not only increases the design cost, but also reduces the possibility of an arbitrary storyline.
@@
The challenge of the drama performance is to redefine the interaction patterns of LLM agents. In many prior studies, these patterns are limited to passive, turn-based responses. For example, some studies focus on the adequacy of single-turn dialogues rather than active participation (Wu et al. 2024). AI agents typically await user instructions, lacking genuine proactivity. This reactive model is hard to support dynamic scenes where multiple AI actors interact with human players. We believe that truly vibrant, dramatic interactions require AI actors to make autonomous decisions, collaborate or conflict in open scenarios, and actively guide the development of the plot. This paradigm shift from passive response to active guidance is the core advocated by Agentic AI (Sapkota, Roumeliotis, and Karkee 2025), providing a new perspective on how to enhance actor initiative in dramatic performance.
@@
In addition to the interaction of agents, a comprehensive framework for drama performance should also incorporate two other critical aspects: physical environment interaction and an overall evaluation system. First, drama is an art of both language and action. Actors’ behaviors should affect their physical environment, and the environment’s feedback is a vital part of the performance. Lacking a physical setting, the drama feels more like an abstract dialogue than a tangible, immersive world. Second, there is currently no effective evaluation method to measure the quality of the drama online performance. Most existing LLM benchmarks focus on text generation quality (Gómez-Rodríguez and Williams 2023) or role-playing ability (Tu et al. 2024; Wu et al. 2025a; Wang et al. 2025), rather than the comprehensive effect of the entire drama performance.
@@
To address the above challenges, we propose HAMLET, a multi-agent framework that enables Hyperadaptive Agent-based Modeling for Live Embodied Theatrics. Specifically, we make the following contributions:
@@
  • A HAMLET framework: Our framework is designed as a multi-agent collaborative workflow, divided into two key stages: offline planning and online performance. Offline planning only needs a simple topic to generate a structured narrative blueprint. This blueprint provides freedom for improvisation of online performance while ensuring the integrity of the main story structure. To enhance the online performance, a Perceive And Decide (PAD) module acts as a pre-response module for each agent, making their perceptions and decisions more human-like.
@@
  • A comprehensive evaluation method and leaderboard: To objectively assess the quality of drama generation and performance, we establish a comprehensive evaluation method that assesses three key dimensions – character performance, narrative quality, and interaction experience. This method incorporates a leaderboard that uses GPT-4o as a strong baseline for win rate comparisons. Additionally, we trained HAMLETJudge, a critic model designed for cost-effective and reliable drama performance evaluation.
@@
  • Extensive experiments: We conducted evaluations of HAMLET’s core components. Notably, both PAD and HAMLETJudge, are 8B models, yet they achieved state-of-the-art performance in their respective tasks.
Related Work
LLM-Based Drama
@@
Research on LLM-based drama has evolved along two main directions: drama generation and drama performance. The initial efforts in drama generation adapted techniques from general story creation, using hierarchical models to plan plots and generate coherent narratives (Fan, Lewis, and Dauphin 2018; Yao et al. 2019). With the development of this field, researchers have tried multi-LLM collaboration (Venkatraman, Tripto, and Lee 2024) and hierarchical methods (Mirowski et al. 2023) that separate planning from generation. A key limitation of these systems is that they require a complete story (Wu et al. 2024) or a leading paragraph (Wu et al. 2025c) as input. Unlike previous work, our work presents a system designed for arbitrary, on-demand drama generation. At the same time, the research focus has also expanded to the second dimension: drama performance. Role-playing in drama performance is usually achieved by LLM agents (Chen et al. 2024b). These agents require knowledge about the actor to learn its expression style (Shanahan, McDonell, and Reynolds 2023; Wang et al. 2023; Tu et al. 2024). Current role-playing systems are built using either prompt engineering (Li et al. 2023; Chen et al. 2023) or fine-tuning (Shao et al. 2023; Lu et al. 2024). However, these methods find it hard to capture complex character traits from simple descriptions (Huang et al. 2023). To create more natural actors, better models are needed. We propose a perceive and decide model to improve dialogue consistency and emotional expression.
Evaluation for Role-Playing Conversation Agents
@@
Traditional dialogue metrics like Embedding Similarity, BLEU, and ROUGE (Mou et al. 2016; Wu et al. 2020; Serban et al. 2017) are widely used in role-playing conversation agent (RPCA) research (Wang et al. 2024; Zhou et al. 2024; Tao et al. 2024). However, these quantitative indicators struggle to evaluate role consistency and personality. To address this, specialised evaluation frameworks have been proposed. RoleEval (Shen et al. 2023) uses role-specific multiple-choice questions to test the model’s understanding of a character. SocialBench (Chen et al. 2024a) constructs evaluation questions from multi-source dialogues, while CharacterEval (Tu et al. 2024) adopts multi-round dialogues and multi-dimensional scoring for aspects like conversational ability. Additionally, RAIDEN (Wu et al. 2025b) builds a Q&A dataset via annotator interaction to evaluate responsiveness in specific dimensions. Although these methods provide quantitative criteria, they are typically limited to two-character or user-character scenarios. While CoSER (Wang et al. 2025) expands the number of roles, it still lacks an evaluation mechanism for overall dramatic performance. Building on this previous work, our study proposes new dimensions to evaluate the complete dramatic performance.
Figure 2: An illustration of HAMLET’s core components for performance generation: a narrative blueprint that defines the scene, plot and character profiles, and the resulting real-time conversation containing scene descriptions and dialogue.
HAMLET: A Multi-Agent Framework Design for AI Drama
@@
To achieve an automated, interactive, and expressive AI drama experience, we designed the HAMLET framework, as shown in Figure 1. The framework decouples the generation and performance of the drama into two main stages: offline planning, a multi-agent collaborative workflow to generate a narrative blueprint, and online performance, a hierarchical control system that executes the blueprint, manages real-time interactions, and handles environmental feedback.
Offline Planning
@@
The target of the offline planning stage is to transform user input into a structured narrative blueprint. This stage is designed to handle two types of input: 1) An arbitrary and customizable topic. For this input type, the offline workflow directly generates the content for a complete act based on the provided topic. 2) A complete literary work. When given a full literary text, the workflow first deconstructs it into a series of acts based on its chapters and content structure. Subsequently, it proceeds with drama design for each act, ensuring the adaptation remains faithful to the source material.
Agent-Based Workflow Architecture
@@
Our offline workflow consists of four agents: the actor designer, the plot designer, the reviewer, and the director. This workflow implements character creation, plot generation, review, and final integration.
Character Profile Generation and Review
@@
The starting point of the workflow is actor construction. The actor designer is responsible for generating actor profiles of core characters based on user input. To ensure the depth and accuracy of character creation, this agent queries an external knowledge base via its search module. The output is a structured actor profile that defines the character’s static attributes, such as background and personality, and dynamic attributes, such as their initial goal and core relationships within the story. The profile is then submitted to the reviewer, who checks the rationality of the character settings, the clarity of motivations, and the relationships between actors.
Plot Generation and Structuring
@@
After all actor profiles are approved, the plot designer composes a preliminary narrative draft based on the topic and actors. This draft is then submitted to the reviewer for evaluation. Once the review is passed, the director is responsible for the final structural processing, reconstructing the linear story draft into a hierarchical plot profile. This process includes the following key steps: 1) Act and scene definition: Divide the drama into several acts and specify the scenes in which each act takes place. 2) Creation of environmental elements: Generate a list of interactive props for each scene. This list has specific descriptions and location information for the props. 3) Definition of points: In each act, define a series of narrative points. Each point contains a clear flag and a result that marks its completion. 4) Backward planning: In order to prevent the plot from deviating, the director will give priority to generating the end point when planning an act. Then, based on this end point, it will supplement and construct the logically coherent preceding points leading to the ending in a reverse manner.

### v2

# Sizhou Chen2,3,4Shufan Jiang1,4Chi Zhang3Xiao-Lei Zhang3Xuelong Li3

@@ Abstract
Creating an immersive and interactive theatrical experience is a long-term goal in the field of interactive narrative. The emergence of large language model (LLM) is providing a new path to achieve this goal. However, existing LLM-based drama generation methods often result in agents that lack initiative and cannot interact with the physical scene. Furthermore, these methods typically require detailed user input to drive the drama. These limitations reduce the interactivity and immersion of online real-time performance. To address the above challenges, we propose 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a multi-agent framework focused on drama creation and online performance. Given a simple topic, the framework generates a narrative blueprint, guiding the subsequent improvisational performance. During the online performance, each actor is given an autonomous mind. This means that actors can make independent decisions based on their own background, goals, and emotional state. In addition to conversations with other actors, their decisions can also change the state of scene props through actions such as opening a letter or picking up a weapon. The change is then broadcast to other related actors, updating what they know and care about, which in turn influences their next action. To evaluate the quality of drama performance generated by 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, we designed an evaluation method to assess three primary aspects, including character performance, narrative quality, and interaction experience. The experimental evaluation shows that 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: can create expressive and coherent theatrical experiences.
1 Introduction
@@
In recent years, large language models (LLMs) have demonstrated strong ability in various generative tasks, especially in creative fields such as story creation (Li et al., 2024) and role-playing (Tu et al., 2024). These models are able to generate fluent and imaginative texts, providing a foundation for the development of interactive narrative. When applying these capabilities to the creation of drama, the exploration can be divided into two parts: drama generation and drama performance.
@@
Existing studies have explored drama generation using language models, employing methods such as hierarchical approaches (Fan et al., 2018; Yao et al., 2019) and the modular generation of script structures (Mirowski et al., 2023). However, significant challenges exist in applying LLMs to structured drama generation, especially in scenarios requiring online real-time enactment. In interactive drama performance, LLM agents take on the role of performing the narrative. In contrast to free-form story creation, drama requires each actor to make decisions and take actions consistent with their profiles and the overall plot progression. These interactions must unfold across a series of scenes to collaboratively advance the narrative. Furthermore, existing methods typically require detailed user input, such as a full story outline (Wu et al., 2024) or elaborate guiding paragraphs (Wu et al., 2025b), which not only increases the design cost, but also reduces the possibility of an arbitrary storyline.
@@
The challenge of the drama performance is to redefine the interaction patterns of LLM agents. In many prior studies, these patterns are limited to passive, turn-based responses. For example, some studies focus on the adequacy of single-turn dialogues rather than active participation (Wu et al., 2024). The agents typically await user instructions, lacking genuine proactivity. This reactive model is hard to support dynamic scenes where multiple AI actors interact with human players. We believe that truly vibrant, dramatic interactions require AI actors to make autonomous decisions, collaborate or conflict in open scenarios, and actively guide the development of the plot. This paradigm shift from passive response to active guidance is the core advocated by Agentic AI (Sapkota et al., 2025), providing a new perspective on how to enhance actor initiative in dramatic performance.
@@
In addition to the intelligent and autonomous interaction of agents, a comprehensive framework for drama performance should also incorporate two other critical aspects: physical environment interaction and an overall evaluation system. First, drama is an art of both language and action. Actors’ behaviors should affect their physical environment, and the environment’s feedback is a vital part of the performance. Lacking a physical setting, the drama feels more like an abstract dialogue than a tangible, immersive world. Second, there is currently no effective evaluation method to measure the quality of realtime drama. Most existing LLM benchmarks focus on text generation quality (Gómez-Rodríguez & Williams, 2023) or role-playing ability (Tu et al., 2024; Wu et al., 2025a; Wang et al., 2025), rather than the comprehensive effect of the entire drama performance.
@@
To address the above challenges, we propose 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a multi-agent framework that enables Hyperadaptive Agent-based Modeling for Live Embodied Theatrics. Specifically, we make the following contributions:
  1. 1.
@@
    A 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: framework: Our framework is designed as a multi-agent collaborative workflow, divided into two key stages: offline planning and online performance. Offline planning only needs a simple topic to generate a structured narrative blueprint. This blueprint provides freedom for improvisation of online performance while ensuring the integrity of the main story structure. To enhance the online performance, a Perceive And Decide (0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:) module acts as a pre-response module for each agent, making their perceptions and decisions more human-like.
  2. 2.
@@
    A comprehensive evaluation method and leaderboard: To objectively assess the quality of drama generation and performance, we establish a comprehensive evaluation method that assesses three curated dimensions. This method incorporates a leaderboard that uses GPT-4o as a strong baseline for win rate comparisons. Additionally, we trained 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a critic model designed for cost-effective and reliable drama performance evaluation.
  3. 3.
@@
    Extensive experiments: We conducted evaluations of 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:’s core components. Notably, both 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, are 8B models, yet they achieved state-of-the-art performance in their tasks respectively.
2 Related Work
@@
LLM-Based Drama. Research on LLM-based drama has evolved along two main directions: drama generation and drama performance. The initial efforts in drama generation adapted techniques from general story creation, using hierarchical models to plan plots and generate coherent narratives (Fan et al., 2018; Yao et al., 2019). With the development of this field, researchers have tried multi-LLM collaboration (Venkatraman et al., 2024) and hierarchical methods (Mirowski et al., 2023) that separate planning from generation. A key limitation of these systems is that they require a complete story (Wu et al., 2024) or a leading paragraph (Wu et al., 2025b) as input. Unlike previous work, our work presents a system designed for arbitrary and customizable drama generation. At the same time, the research focus has also expanded to the second dimension: drama performance. Role-playing in drama performance is usually achieved by LLM-based agents (Chen et al., 2024b). These agents require knowledge about the actor to learn its expression style (Shanahan et al., 2023; Wang et al., 2023; Tu et al., 2024). Current role-playing systems are built using either prompt engineering (Li et al., 2023; Chen et al., 2023) or fine-tuning (Shao et al., 2023; Lu et al., 2024). However, these methods find it hard to capture complex character traits from simple descriptions (Huang et al., 2023). To create more natural actors, better models are needed. We propose a perceive and decide model to improve dialogue consistency and emotional expression.
@@
Evaluation for Role-Playing Conversation Agents. Traditional dialogue metrics like Embedding Similarity, BLEU, and ROUGE (Mou et al., 2016; Wu et al., 2020; Serban et al., 2017) are widely used in role-playing conversation agent (RPCA) research (Wang et al., 2024; Zhou et al., 2024; Tao et al., 2024). However, these quantitative indicators struggle to evaluate role consistency and personality. To address this, specialised evaluation frameworks have been proposed. RoleEval (Shen et al., 2023) uses role-specific multiple-choice questions to test the model’s understanding of a character. SocialBench (Chen et al., 2024a) constructs evaluation questions from multi-source dialogues, while CharacterEval (Tu et al., 2024) adopts multi-round dialogues and multi-dimensional scoring for aspects like conversational ability. Additionally, RAIDEN (Wu et al., 2025a) builds a Q&A dataset via annotator interaction to evaluate responsiveness in specific dimensions. Although these methods provide quantitative criteria, they are typically limited to two-character or user-character scenarios. While CoSER (Wang et al., 2025) expands the number of roles, it still lacks an evaluation mechanism for overall dramatic performance. Building on this previous work, our study proposes new criterias to evaluate the complete dramatic performance.
3 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:: A Multi-Agent Framework for AI Drama
@@
To achieve an automated, interactive, and expressive AI drama experience, we designed the 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: framework, as shown in Figure 1. The framework decouples the generation and performance of the drama into two main stages: offline planning, a multi-agent collaborative workflow to generate a narrative blueprint, and online performance, a hierarchical control system that executes the blueprint, manages real-time interactions, and handles environmental feedback.
Figure 1: The 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: framework creates AI drama in two main stages. First, during offline planning, a collaborative workflow of agents including the actor designer, plot designer, and reviewer creates initial materials, which are then integrated by a director agent into a structured narrative blueprint. This blueprint then guides the subsequent online performance, where a control system composed of a planner, transfer, and advancer directs a dynamic and improvisational theatrical experience. The figure illustrates this entire process using the play-within-a-play scene from Hamlet, Act III, Scene II.
3.1 Offline Planning
@@
The target of the offline planning stage is to transform user input into a structured narrative blueprint. This stage is designed to handle two types of input: i) An arbitrary and customizable topic. For this input type, the offline workflow is able to generate a complete act based on the provided topic. ii) A complete literary work. When given a full literary text, the workflow first deconstructs it into a series of acts based on its chapters and content structure. Subsequently, it proceeds with drama design for each act, ensuring the adaptation remains faithful to the source material.
@@
Agent-Based Workflow Architecture. The offline workflow consists of four agents: the actor designer, the plot designer, the reviewer, and the director. This workflow implements character creation, plot generation, review, and final integration.
@@
Character Profile Generation and Review. The starting point of the workflow is actor construction. The actor designer is responsible for generating actor profiles of core characters based on user input. To ensure the depth and accuracy of character creation, this agent queries an external knowledge base via its search module. The output is a structured actor profile that defines the character’s static attributes, such as background and personality, and dynamic attributes, such as their initial goal and core relationships within the story. The profile is then submitted to the reviewer, who checks the rationality of the character settings, the clarity of motivations, and the relationships between actors.
Figure 2: An illustration of 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:’s core components for performance generation: a narrative blueprint that defines the scene, plot and character profiles, and the resulting real-time conversation containing scene descriptions and dialogue.
@@
Plot Generation and Structuring. After all actor profiles are approved, the plot designer composes a preliminary narrative draft based on the topic and actors. This draft is then submitted to the reviewer for evaluation. Once the review is passed, the director is responsible for the final structural processing, reconstructing the linear story draft into a hierarchical plot profile. This process includes the following key steps: 1) Act and scene definition: Divide the drama into several acts and specify the scenes in which each act takes place. 2) Creation of environmental elements: Generate a list of interactive props for each scene. This list has specific descriptions and location information for the props. 3) Definition of points: In each act, define a series of narrative points. Each point contains a clear flag and a result that marks its completion. 4) Backward planning: In order to prevent the plot from deviating, the director will give priority to generating the end point when planning an act. Then, based on this end point, it will supplement and construct the logically coherent preceding points leading to the ending in a reverse manner.